# Navigating the Landscape of Hint Generation Research: From the Past to the Future

### v1

@@ 4.2 Components of an Effective Hint Generation System
In this section, we briefly review some research areas that can inform the design of a hint generation system that aligns with the formal definition.
Question Answering. Current hint generation systems are framed as scaffolding tools within a question answering (QA) setup, where learners are assessed on their ability to answer questions on a relevant topic. These systems can benefit from research in question answering in three ways - firstly, by equipping the hint generation models with the reasoning abilities to answer questions. Anderson et al. (1995) emphasizes on the significance of an underlying production-rule model that helps to break down a learning goal into achievable subgoals in their cognitive tutoring framework, and we posit the QA systems today can form the foundation of this production-rule model. Datasets like StrategyQA Geva et al. (2021) and strategies like Self-Ask Press et al. (2022) and Socratic Questioning Qi et al. (2023) are great illustrations of the goal decomposition ability described in the ACT-R cognitive learning theory framework Anderson et al. (1995). Secondly, question answering systems are quite diverse, and can solve complex questions spread across multiple modalities like knowledge bases LAN et al. (2022), tables Jin et al. (2022), images Srivastava et al. (2021); de Faria et al. (2023), videos Zhong et al. (2022) etc., as well as explain the reasoning for the answer Danilevsky et al. (2020); Schwalbe and Finzel (2023). These question answering models can form the basis of hint generation systems that take into account incomplete solutions, and the necessary reasoning steps to answer the question. Thirdly, QA efforts can help provide seed datasets and data annotation strategies to expand the hint generation resources. The literature is rich in education-related datasets ranging from generic datasets like RACE Lai et al. (2017), LearningQ Chen et al. (2018), TQA Kembhavi et al. (2017) etc. to subject-specific datasets like SciQ Welbl et al. (2017), ScienceQA Lu et al. (2022), SituatedQA Zhang and Choi (2021), GSM8K Cobbe et al. (2021), BioASQ Tsatsaronis et al. (2015), CORD-19 Wang et al. (2020), PubMedQA Jin et al. (2019) etc.. We could extend these datasets, by taking a collaborative approach between the ML practitioners and expert tutors to explore the utilization of hint in classroom or one-on-one tutoring setting across different domains, and develop better hint generation datasets similar to those outlined in Section 3.
@@
Answer Assessment. Answer assessment plays a pivotal role in evaluating and understanding learner responses, enabling the system to provide targeted feedback and adaptive guidance. Accurate answer assessment is foundational for standardizing the grading process, identifying misconceptions, tracking individual progress, and tailoring subsequent hints to address specific learning needs. The literature includes several answer evaluation strategies, varying from naive exact match and token overlap approaches, to BERT-based semantic strategies for short form answers Bulian et al. (2022), explainable systems trained on LLM-distilled rationales Li et al. (2023), semantic grouping based systems for batch grading Chang et al. (2022) and multi-modal assessment frameworks to evaluate oral presentations Liu et al. (2020). For detailed overview of answer assessment strategies, we point the readers to the survey written by Das et al. (2021).
@@
User Modeling. User modeling is extensively explored within the realm of recommendation systems, with the aim to adapt and customize a service to user’s specific needs. User modeling holds the key to identifying the learner’s interests and preferences within the hint generation space, (Equation 4) helping align the hints to learner’s prior knowledge (Section 2.2) and personalize the feedback 2.1. Liu et al. (2022), for instance, utilizes long short-term memory (LSTM) cells Hochreiter and Schmidhuber (1997) for knowledge estimation Corbett and Anderson (1994) of student’s current understanding based on their past responses for open-ended program synthesis. Similar strategies to model student’s progress might help enhance their learning experience, and direct the readers to comprehensive surveys authored by Li et al. (2021) and He et al. (2023) for an extensive overview.
Question Generation. Questions play pivotal roles in education, serving to recall knowledge, test comprehension, and foster critical thinking. Al Faraby et al. (2023) classify neural question generation for educational purposes into three broad categories - Question Generation from Reading Materials, Word Problem Generation, and Conversation Question Generation. We posit a similar use case for question generation to generate hints within conversations, to clarify not just the learner’s response, but also their understanding and ability to present their ideas. We point the readers to surveys written by Kurdi et al. (2020), Zhang et al. (2021), Das et al. (2021) and Pan et al. (2019) for a comprehensive overview.
@@
Dialogue Modeling. Intelligent tutoring systems are mediated by dialogue systems (or conversational agents), providing an interface between ITS and learners. With the recent developments in retrieval augmented generation Gao et al. (2023); Cai et al. (2022); Li et al. (2022), emotion-aware dialogue systems Ma et al. (2020), and enhanced language understanding capabilities in multi-turn dialogues Zhang and Zhao (2021), dialogue systems today can also help improve the hint generation models, empowering them with abilities to better understand the learner’s queries, understand and respond to complex domains, generate affective responses, and keep track of long conversation sessions for adaptive tutoring that would help link the generated hints to learner’s prior knowledge (Section 2.2). For further exploration, the readers are invited to read the surveys written by Ni et al. (2023), Deriu et al. (2021) and Ma et al. (2020).
4.3 Open Challenges for an Effective Hint Generation Systems
@@
In this section, we outline challenges and future directions for building more effective hint generation systems that align with our formal definition.
@@
Privacy-preserving self-evolving frameworks. Current hint generation frameworks limit their applications to a fixed dataset or a pre-defined set of problems, which does not scale well in real-world applications. In order to make these systems more robust and effective, we need them to incorporate the learner’s feedback as well as their implicit preferences (refer to Equations 3 and 4). Thus, it is important to design self-evolving frameworks that provide personalized feedback, while respecting the learner’s rights for their privacy. Further research is needed to incorporate user-modeling and explore privacy-related approaches like differential privacy Zhao and Chen (2022); Yang et al. (2023), federated learning Li et al. (2021) to create adaptive systems ethically.
@@
Diverse domain exploration. Most efforts within the hint generation space are limited to programming, language acquisition or mathematics, similar to the domain trends in intelligent tutoring systems Mousavinasab et al. (2021). However, the realm of education and tutoring extends beyond these fields into other subjects within natural sciences (e.g., physics, chemistry, biology, earth sciences), social sciences (e.g., history, civics, geography, law) and learning beyond educational institutions (e.g., educating patients about their health conditions effectively Gupta et al. (2020)). Tackling the hint generation problem in these subjects raises several challenges such as the evaluation of subjective long-form answers, the need of domain knowledge and enhanced reasoning capabilities Lyu et al. (2024); Huang and Chang (2022); Zhang et al. (2023). However, it simultaneously opens up new opportunities to expand the capabilities of hint generation systems, including development of smaller models that combine the power of pre-trained language models with the power of adapted knowledge models (e.g., COMET Bosselut et al. (2019)).
@@
Multi-lingual and multi-cultural aspects. Prior studies have found a correlation between the linguistic and cultural diversity and capabilities for innovation Hofstra et al. (2020); Evans and Levinson (2009). However, we found the majority of research at the intersection of natural language processing and learning sciences is limited to the English language, either as the mode of education, or a subject for language acquisition. Therefore, there is an urgent need to develop multi-lingual, culturally-aware systems that accommodate the education for non-English speakers, and sustain the diverse cultural and linguistic heritage Bernard (1992).
@@
Multi-modal elements. Prior qualitative research has established gains from complementing education with additional modalities like maps, diagrams, and multi-media content Winn (1991); Swidan and Naftaliev (2019); Tippett (2016); Kapi et al. (2017); Abdulrahaman et al. (2020). We believe incorporating these cues to a hint generation system would enhance the students’ learning experience, improve their understanding, enhance the memory, and overall help create a more holistic tool for the education landscape Pourkamali et al. (2021).
@@
Affective systems. Affective aspects are often neglected when building intelligent tutoring systems and hint generation systems Hasan et al. (2020). However, incorporating them into the education pipeline is key for personality development, encouragement and improving self-motivation Jiménez et al. (2018). Jiménez et al. (2018) shows that using affective feedback has a positive impact on students facing learning challenges. Thus, affective hint generation systems that take into consideration the emotional state of the learners when providing feedback will be an important area of study.
@@
Accessible systems. Another fruitful and challenging area of research is to develop accesible hint generation systems. For people with neurodevelopmental disorders (like attention-deficit/hyperactivity disorder i.e. ADHD and autism spectrum disorder i.e. ASD) and learning disabilities (like dyslexia and dyscalculia), one can modify the hint generation systems to have: a) simplified text: using plain language, and avoid using jargons and complex terminology Štajner (2021), b) multisensory supports: leveraging a combination of visual, auditory, and kinesthetic modalities to present hints in multiple formats Vezzoli et al. (2017); Gori and Facoetti (2014), c) interactive elements: incorporating interactive elements to engage learners to explore concepts in a hands-on manner García-Carrión et al. (2018), and d) predictable routine: establishing a consistent routine for delivering hints consistently can helps learners feel more comfortable and confident Love et al. (2012). Creating a holistic accessible learning system is a shared responsibility, and needs to be further explored to improve lives of people with disabilities.
@@
Evaluation metrics. Evaluation of generated hints is a non-trivial task, depending on multiple factors to determine success of a hint. Most community determines the success of a hint generation system by the learner’s abilities in producing the reference solution (Equation 2). However, this evaluation framework does not penalize the generated hints if they make the problem too simple, and also does not take into account the individual learner’s preferences and learning objectives. Therefore, we need to build human-centered evaluation frameworks Lee et al. (2022) that can help measure factors beyond the learner’s answering capabilities, such as learner’s ownership over the learning process, long-term capabilities of generated knowledge, motivation and enjoyment they received during their interactions and so forth.
@@ 4.4 Ethical considerations
We must collectively explore the broader implications of integrating NLP in education on representativeness and equity and exacerbating systemic inequalities (Weidinger et al., 2022). There is a risk that the hints generated by NLP models may not adequately reflect the diverse backgrounds and lived experiences of students (Dixon-Román et al., 2020) and potentially perpetuate harmful stereotypes about different identities (Dev et al., 2021). Prior work has demonstrated the various forms of ‘bias’ in NLP systems (Blodgett et al., 2020), which may contribute to the construction of language hierarchies and limiting language variation (Schneider, 2022). To promote inclusive design and mitigate these ethical considerations, it is essential to understand how power, privilege, and resources are redistributed as a result of introducing AI in the classroom. Is there a possibility of diminishing quality education for marginalized and under-resourced groups (Remian, 2019)? We must take a community-collaborative approach to understand how to design justice-oriented and accountable systems where learners can truly benefit from hint generation systems (Madaio et al., 2022).

### v2

@@ 4.2 Components of an Effective Automatic Hint Generation System
In this section, we briefly review some NLP research areas that can inform the design of an automatic hint generation system (Figure 3) that aligns with the proposed formal definition introduced in Section 4.1.
Question Answering. Current hint generation systems are framed as scaffolding tools within a question answering (QA) setup, where learners are assessed on their ability to answer questions on a relevant topic. These systems can benefit from research in question answering in two ways.
Firstly, hint generation models should be equipped with reasoning abilities to answer questions. Anderson et al. (1995) emphasizes the significance of an underlying production-rule model that helps to break down a learning goal into achievable subgoals in their cognitive tutoring framework, and we posit the QA systems today can form the foundation of this production-rule model. Datasets like StrategyQA Geva et al. (2021) and strategies like Self-Ask Press et al. (2022) and Socratic Questioning Qi et al. (2023) are great illustrations of the goal decomposition ability described in the ACT-R cognitive learning theory framework Anderson et al. (1995). We can adapt these computational question answering frameworks coupled with the answer assessment module (described below) to learn mapping $\textbf{H}:\{q,a,\mathcal{K}_{q\rightarrow a},D^{l}_{q}\}\rightarrow\mathcal{H}$.
Secondly, question answering systems are quite diverse and can solve complex questions spread across multiple modalities such as knowledge bases LAN et al. (2022), tables Jin et al. (2022), images Srivastava et al. (2021); de Faria et al. (2023), and videos Zhong et al. (2022), as well as explain the reasoning for the answer Danilevsky et al. (2020); Schwalbe and Finzel (2023). These question answering models can form the basis of hint generation systems that take into account incomplete solutions and the necessary reasoning steps to answer the question.
Answer Assessment. Answer assessment plays a crucial role in evaluating and understanding learner responses, enabling the system to provide targeted feedback and adaptive guidance. Accurate answer assessment is foundational for standardizing the grading process, identifying misconceptions, tracking individual progress, and tailoring subsequent hints to address specific learning needs. We can utilize an answer assessment module to identify the missing (or wrong) components of an ongoing answer ($\hat{a}_{k}$) juxtaposed to a reference answer ($a$) and provide appropriate hints to scaffold them towards a correct solution. We can also utilize the answer assessment module to aggregate the mistakes made by a learner over time to track their progress.
@@
Prior work includes several answer evaluation strategies, varying from naive exact match and token overlap approaches to BERT-based semantic strategies for short form answers Bulian et al. (2022), explainable systems trained on LLM-distilled rationales Li et al. (2023a), semantic grouping based systems for batch grading Chang et al. (2022) and multi-modal assessment frameworks to evaluate oral presentations Liu et al. (2020). For a detailed overview of answer assessment strategies, we point the readers to the survey written by Das et al. (2021).
User Modeling. User modeling is extensively explored within the recommendation systems research area, with the aim of adapting and customizing a service to a user’s specific needs. User modeling holds the key to identifying the learner’s interests and preferences within the hint generation framework (Equation 4), helping align the hints to the learner’s prior knowledge (Section 2.2) and personalize the feedback (Section 2.1). Using a learner’s past interaction data ($\mathcal{L}_{l}$), we can learn a preference function ($\mathcal{F}^{l}_{pref}$) based on the success of previous hints in helping the learner achieve their learning goal. We can measure this learning goal ($\mathcal{F}^{l}_{learning}$) using direct or indirect learner feedback from the user interaction. For instance, we can obtain learners’ self-reported data about the effectiveness of a hint, or develop alternate indicators of success by observing the learner’s behavior across several dimensions, such as the number of interactions, response initiation time, degree of change in response, and frequency of hints asked.
@@
Liu et al. (2022), for instance, utilizes long short-term memory (LSTM) cells Hochreiter and Schmidhuber (1997) for knowledge estimation Corbett and Anderson (1994) of student’s current understanding based on their past responses for open-ended program synthesis. They utilize these time-varying student knowledge states to predict students’ responses to programming problems, monitoring and analyzing their progress. We can adopt similar user modeling techniques to measure progress, adapt the hints to the learner’s preferences, and design a curriculum (or suggest areas) for further improvement. We direct the readers to the comprehensive survey authored by He et al. (2023) for an extensive overview on user modeling.
@@
Dialogue Generation. Intelligent tutoring systems are mediated by dialogue systems (or conversational agents), providing an interface between ITS and learners. Recent developments in retrieval augmented generation Gao et al. (2023); Cai et al. (2022); Li et al. (2022), emotion-aware dialogue systems Ma et al. (2020), and enhanced language understanding capabilities in multi-turn dialogues Zhang and Zhao (2021) can help improve the hint generation models, empowering them with abilities to better understand the learner’s queries, understand and respond to complex domains, generate affective responses, and keep track of long conversation sessions for adaptive tutoring that would help link the generated hints to learner’s prior knowledge (Section 2.2). For further exploration, we invite readers to read the surveys written by Ni et al. (2023), Deriu et al. (2021) and Ma et al. (2020).
Question Generation. Questions play a pivotal role in education, serving to help recall knowledge, test comprehension, and foster critical thinking. Al Faraby et al. (2023) classify neural question generation for educational purposes into three broad categories: (1) Question Generation from Reading Materials, (2) Word Problem Generation, and (3) Conversation Question Generation. We posit a similar use case for question generation to generate hints within conversations to clarify not just the learner’s response but also their understanding and ability to present their ideas. We point the readers to surveys written by Kurdi et al. (2020), Zhang et al. (2021), Das et al. (2021) and Pan et al. (2019) for a comprehensive overview.
Modular structure of proposed approach. Although an ideal hint generation system would be able to satisfy all the criterion described in our formal definition (Equations 1, 2, 3 and 4)—providing hints that are personalized to individual learners preferences and learning objectives—we acknowledge the limitations of technology at the time of this survey (Section 4.3). Keeping that in mind, our proposed formal definition (Section 4.1) and components of hint generation system (Section 4.2) suggests a modular structure, comprising of some essential components (such as Equations 1 and 2 that form the crux of a hint generation system, or the question answering module that helps develop the reasoning to develop the hints) and some complementary modules (like Equations 3 and 4 that helps personalize the hints to individual learners or question generation module that adapts particular hint giving strategies) to further enhance the quality of generated hints. This paper proposes thus a long term roadmap for automatic hint generation systems that can showcase intermediate progress along the way.
4.3 Open Challenges for Effective Automatic Hint Generation Systems
@@
In this section, we outline challenges and future directions for building more effective automatic hint generation systems that align with our formal definition.
@@
Privacy-preserving self-evolving frameworks. Current hint generation frameworks limit their applications to a fixed dataset or a pre-defined set of problems, which does not guarantee high quality performance in real-world applications. In order to make these frameworks more effective for the learners, we need to adopt self-evolving frameworks that can incorporate the learner’s feedback and implicit preferences. These self-evolving frameworks should have the capability of identifying user feedback from an ongoing conversation ($D^{l}_{q}$), gathering implicit preferences ($\mathcal{F}^{l}_{pref}$) from past learning interactions ($\mathcal{L}_{l}$), and adapting their hint generation process to an individual learner while respecting their right to privacy.
Although there is prior work on incorporating human-feedback to improve generation quality (e.g., prompt optimization Chang et al. (2024); Sahoo et al. (2024), reinforcement learning with human feedback Kaufmann et al. (2023); Casper et al. (2023)), incorporating these user-modeling aspects in a privacy-preserving manner remains an active area of research Miranda et al. (2024). Differential privacy Zhao and Chen (2022); Yang et al. (2023) and federated learning Li et al. (2021) could be relevant research avenues to help create these self-evolving frameworks with a responsible approach. For an in-depth review of the state of preserving privacy in LLMs, readers are encouraged to consult the survey by Miranda et al. (2024).
@@
Diverse domain exploration. Most efforts within the hint generation space are limited to programming, language acquisition or mathematics, similar to the domain trends in intelligent tutoring systems Mousavinasab et al. (2021). However, education and tutoring often involve other subjects within natural sciences (e.g., physics, chemistry, biology, earth sciences), social sciences (e.g., history, civics, geography, law), and learning beyond educational institutions (e.g., educating patients about their health conditions effectively Gupta et al. (2020)). Tackling the hint generation problem in these subjects raises several challenges, such as the evaluation of subjective long-form answers, the need for domain knowledge and enhanced reasoning capabilities Lyu et al. (2024); Huang and Chang (2022); Zhang et al. (2023). However, it simultaneously opens up new opportunities to expand the capabilities of hint generation systems, including the development of smaller models that combine the power of pre-trained language models with the power of adapted knowledge models (e.g., COMET Bosselut et al. (2019)).
Efforts in question answering can also help provide seed datasets and data annotation strategies to expand the hint generation resources. The literature is rich in education-related datasets ranging from generic datasets like RACE Lai et al. (2017), LearningQ Chen et al. (2018), TQA Kembhavi et al. (2017) etc. to subject-specific datasets like SciQ Welbl et al. (2017), ScienceQA Lu et al. (2022), SituatedQA Zhang and Choi (2021), GSM8K Cobbe et al. (2021), BioASQ Tsatsaronis et al. (2015), CORD-19 Wang et al. (2020), PubMedQA Jin et al. (2019). To extend these datasets we envision a collaborative approach between the ML practitioners and expert tutors to explore the utilization of hints in classroom or one-on-one tutoring setting across different domains.
@@
Multi-lingual and multi-cultural aspects. Prior studies have found a correlation between linguistic and cultural diversity and capabilities for innovation Hofstra et al. (2020); Evans and Levinson (2009). However, we found that the majority of research at the intersection of natural language processing and learning sciences is limited to the English language, either as the mode of education or a subject for language acquisition. Although utilizing multi-lingual large language models (mLLMs) for hint generation can potentially help incorporate the linguistic aspects of the learning experience, providing culturally-aware hints still remains a challenging task. Liu et al. (2024) reveals that while mLLMs are aware of cultural proverbs, they struggle to reason with figurative proverbs and sayings. Building benchmarks to evaluate cultural awareness of generation models such as MAPS Liu et al. (2024), PARIKSHA Liu et al. (2024), CultureAtlas Fung et al. (2024) and CUNIT Li et al. (2024) is an essential first step towards identifying the shortcomings of hint generation frameworks. Moving forward, we need to explore the hint generation capabilities of mLLMs and develop culturally-aware systems that accommodate the education for non-English learners, and sustain the diverse cultural and linguistic heritage Bernard (1992); Soto et al. (1999).
@@
Multi-modal elements. Prior qualitative research has established gains from complementing education with additional modalities like maps, diagrams, and multi-media content Winn (1991); Swidan and Naftaliev (2019); Tippett (2016); Kapi et al. (2017); Abdulrahaman et al. (2020). Research in intelligent tutoring systems has also explored incorporating certain gamifying elements like badges, leaderboards, narratives, and virtual currency to keep the learners more engaged and motivated González et al. (2014); Ramadhan et al. (2024). We believe incorporating these cues into a hint generation system would improve the students’ memory, understanding, and overall learning experience to help create a holistic educational tool Pourkamali et al. (2021).
@@
Affective systems. Affective aspects are often neglected when building intelligent tutoring systems and hint generation systems Hasan et al. (2020). However, incorporating them into the education pipeline is key for personality development, encouragement and improving self-motivation Jiménez et al. (2018). Jiménez et al. (2018) shows that using affective feedback has a positive impact on students facing learning challenges. Thus, affective hint generation systems that take into consideration the emotional state of the learners when providing feedback is an important area of study.
In recent years, there has been increased interest in building dialogue agents that can adapt to the emotional state of the user in an ongoing dialogue setting. Haydarov et al. (2023) develops a large-scale multi-modal benchmark for visually grounded emotional reasoning-based conversations that use visual stimuli to stir a conversation to test out the emotional reasoning capabilities of multi-modal systems. Li et al. (2023b), on the other hand, proposed a future emotion state prediction framework in spoken dialogue systems to predict the future affective reactions of users based on the ongoing conversation. Resources and findings from the ongoing research in affective dialogue systems could be leveraged and advanced to develop more adaptive and emotion-aware tutoring systems. We point the readers to Ma et al. (2020), Raamkumar and Yang (2022), and Zhang et al. (2024) for a comprehensive review of affective dialogue systems.
@@
Accessible systems. Another fruitful and challenging area of research is to develop accesible hint generation systems. For people with neurodevelopmental disorders (e.g., attention-deficit/hyperactivity disorder (ADHD) and autism spectrum disorder (ASD)) and learning disabilities (e.g., dyslexia and dyscalculia), one can modify the hint generation systems to have: (1) simplified text: using plain language, and avoid using jargons and complex terminology Štajner (2021), (2) multi-sensory supports: leveraging a combination of visual, auditory, and kinesthetic modalities to present hints in multiple formats Vezzoli et al. (2017); Gori and Facoetti (2014), (3) interactive elements: incorporating interactive elements to engage learners to explore concepts in a hands-on manner García-Carrión et al. (2018), and (4) predictable routine: establishing a consistent routine for delivering hints consistently can helps learners feel more comfortable and confident Love et al. (2012).
@@
Evaluation metrics. Evaluation of generated hints is a non-trivial task, depending on multiple factors to determine the quality and the success of a hint. Prior work determines the success of a hint generation system by the learner’s abilities in producing the reference solution (Equation 2) Jatowt et al. (2023). However, this evaluation framework does not penalize the generated hints if they make the problem too simple, and also does not take into account the individual learner’s preferences and learning objectives. Therefore, we need to build human-centered evaluation frameworks Lee et al. (2022) that can help measure factors beyond the learner’s answering capabilities, such as learner’s ownership over the learning process, long-term capabilities of generated knowledge, motivation and enjoyment they receive during their interactions.
@@ 4.4 Ethical considerations
We must collectively explore the broader implications of integrating NLP in education on representativeness and equity and exacerbating systemic inequalities (Weidinger et al., 2022). There is a risk that the hints generated by NLP models may not adequately reflect the diverse backgrounds and lived experiences of students (Dixon-Román et al., 2020) and potentially perpetuate harmful stereotypes about different identities (Dev et al., 2021). Prior work has demonstrated the various forms of ‘bias’ in NLP systems (Blodgett et al., 2020), which may contribute to the construction of language hierarchies and limiting language variation (Schneider, 2022). To promote inclusive design and mitigate these ethical considerations, it is essential to understand how power, privilege, and resources are redistributed as a result of introducing AI in the classroom. Is there a possibility of diminishing quality education for marginalized and under-resourced groups (Remian, 2019)? We must take a community-collaborative approach to understand how to design justice-oriented and accountable systems (Madaio et al., 2022) where learners can truly benefit from hint generation systems.